Find the equation of the curve which passes through the point (1 , 4) and is such that : dy÷dx =2x^2 +3x +2
1
Expert's answer
2019-09-20T11:04:27-0400

Let us rewrite the differential equation in form "dy = (2 x^2 + 3 x + 2 ) dx", and integrate both sides: "y = \int dy = \int (2 x^2 + 3 x + 2 ) dx =\frac{2}{3} x^3 + \frac{3}{2} x^2 + 2 x + C".

Substituting the boundary condition "y(1) = 4" into last equation, obtain "4 = \frac{25}{6} + C", from where "C = -\frac{1}{6}". Hence, the solution of given differential equation is "y(x) = \frac{2}{3}x^3 + \frac{3}{2} x^2 + 2 x - \frac{1}{6}" .